OpenStudy (anonymous):

medal please help Use the table below to answer this question: x1 y10 x3 y4 x5 x-2 Find the average rate of change for the given function from x = 3 to x = 5. −6 −3 3 6

OpenStudy (anonymous):

would it be -6

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

thx again

OpenStudy (anonymous):

tj

OpenStudy (anonymous):

no rush

OpenStudy (texaschic101):

average rate of change is the slope slope = (y2 - y1) / (x2 - x1) (1,10)....x1 = 1 and y1 = 10 (3,4)...x2 = 3 and y2 = 4 now we sub slope = (4 - 10) / (3 - 1) slope = -6/2 slope = -3 you might want to get a second opinion on this....I am not too sure about this

OpenStudy (anonymous):

ohok